Connect Homework - Chapter 9 A Saved Help Save & Exit Submit Sales Variable expenses (60%) Income before interest Interest ex I need help with the solution and explanation in the whole question Please and Thank you.
Answer #1

You can calculate this easily using operating leverage method.

Degree of operating leverage (DOL) = Income before interest/Net income

DOL for Miller = 230,000/154,000 = 1.49

DOL for Weaver = 460,000/154,000 = 2.99

The formula now is

Increase in net income = DOL * Increase in sales

Miller (1.49*20)% increases by 30%
Weaver (2.99*20)% increases by 60%

Note :

I have rounded some numbers

Comment if you face any issues